OpenStudy (anonymous):

1-Using the words: odd; even; or neither, classify each of the following functions.U must provide supporting evidence for your classification? 1- f(x) - 2x + 1 2- f(x) 4x^2 - 3 3- f(x) 3x^3 - 2x^2 + 5 4- f(x) 1/x^3 -5x

OpenStudy (welshfella):

for a function to be even f(x) = f(-x) for an odd function f(-x) = - f(x)

OpenStudy (welshfella):

so for Number 2 f(1) = 4(1)^2 -3 = 4 -3 = 1 f(-1) = 4(-1)^2 - 3 = 4 - 3 = 1 so this is an even function

OpenStudy (welshfella):

even functions are symmetrical about the y-axis

OpenStudy (anonymous):

What about the answer of 3 and 4 can you also help in it plz

OpenStudy (welshfella):

and odd functions have got origin symmetry

OpenStudy (welshfella):

find f(1) and f(-1) for number 3 what do you get?

OpenStudy (welshfella):

3(1)^3 - 2((1)^2 + 5 = ? 3(-1)^3 - 2((-1)^2 + 5 = ?

OpenStudy (welshfella):

if it does ot fit the requirements for even and odd then it is neither

OpenStudy (welshfella):

3(1)^3 - 2((1)^2 + 5 = 3 - 2 + 5 = 6 3(-1)^3 - 2((-1)^2 + 5 = -3 -2 + 5 = 0

OpenStudy (welshfella):

- so that is neither

OpenStudy (welshfella):

sorry gotta go right now

OpenStudy (anonymous):

and what abut fourth one ?

OpenStudy (anonymous):

thnak you for your help i ll try 4th by my self you are genius :)

OpenStudy (anonymous):

thanks

Vocaloid (vocaloid):

it's the same process as welsh said for a function to be even f(x) = f(-x) for an odd function f(-x) = - f(x) to see whether f(x) = 1/x^3 -5x is even or odd, we can use two steps: 1. to test whether f(x) is even, evaluate f(1) and f(-1). if they are the same, then the function is even. otherwise, the function is not even. 2. to test whether f(x) is odd, evaluate f(-1) and -f(1), if they are the same, then the function is odd. otherwise, the function is not odd

OpenStudy (anonymous):

so can you help in 4th one ?